# On the torsion locus of the Ceresa normal function

### v1

@@ 1. Introduction
Let $g\geq 1$ and let $C$ be a smooth complex projective curve of genus $g$. A point $x\in C(\mathbb{C})$ defines the Abel-Jacobi embedding $C\hookrightarrow J(C)$, where $J(C)$ is the Jacobian of $C$. Let $\iota\colon J(C)\rightarrow J(C)$ be the involution $y\mapsto-y$. Then the cycle $Z_{C}:=[C]-[\iota(C)]\in\mathrm{CH}^{g-1}(J(C))$ is a homologically trivial cycle on $J(C)$, and has a well defined image in $\mathrm{Griff}_{1}(J(C))$, the group of homologically trivial cycles modulo the group of algebraically trivial cycles (of dimension one). A famous theorem of Ceresa [Cer83] states that for $g\geq 3$, a very general smooth projective curve $C$ of genus $g$ has non-torsion Ceresa cycle in $\mathrm{Griff}_{1}(J(C))$.
@@
Let $g\geq 3$ and let $H_{g}$ denote the hyperelliptic locus in $\mathcal{M}_{g}$. There exists a finite set of subvarieties $T_{i}$ of dimension $\leq 2g-2$, and a finite set of preimages of Shimura subvarieties $X_{j}$, such that if $U$ is the complement of $(\bigcup_{i}T_{i})\cup(\bigcup_{j}X_{j})\cup H_{g}$ in $\mathcal{M}_{g}$, then the locus in $U(\mathbb{C})$ where $\nu$ is torsion is a countable set of points. Moreover, any positive dimensional component of the torsion locus of $\nu$ with Mumford-Tate group $\mathrm{GSp}_{2g}$ is defined over $\overline{\mathbb{Q}}$.
@@
The locus where $Z_{C}$ itself is torsion is contained in the torsion locus of $\nu$. In particular, by the standard spreading-out argument, the torsion locus of $Z_{C}$ in $U$ is a countable subset of $U(\overline{\mathbb{Q}})$. We note that a similar result has been recently obtained by S-W. Zhang.
@@ 1.1. Ideas of the proofs
The proofs of the previous theorems rely on arithmetic and functional transcendence results in Hodge theory, namely the Manin–Mumford conjecture, Ax–Schanuel Theorem for variations of mixed Hodge structures recently established by Gao–Klingler [GK24], and recent results on the study of the Hodge locus for variations of mixed Hodge structures proved by Baldi–Urbanik [BU24]. The equidistribution part in Theorem 1.5 relies on the computation of the generic rank of normal functions with big monodromy and general equidistribution results established in [TT23].
@@ 1.2. Acknowledgement
The present work started during the conference “The Ceresa Cycle in Arithmetic and Geometry” at ICERM. We thank the organizers for their efforts and ICERM for its hospitality, and Philip Griffiths and Jef Laga for valuable discussions. M.K. was supported by NSF grant DMS-2101482, and S.T. was supported by NSF grant DMS-2302388.
@@ 3.2. Analytic density and equidistribution of the torsion locus
The following theorem is inspired by [KU23]; see also [Gao20] for earlier results.
@@ 4.1. The torsion locus for the Ceresa cycle
We know from [LS23] that in genus 3, there is a ball quotient $Y\subset\mathcal{A}_{3}$ ruled by rational curves, countably many of which are contained in $\Phi(\mathrm{TL}_{\text{pos}}(\nu))$ for the Ceresa normal function. The torsion orders here are unbounded and the torsion locus is Zariski dense, as predicted by Theorem 3.9 and Remark 3.10 (as $\nu|_{Y}$ takes values in the Jacobian of an isotrivial VHS of rank 2); while the fact that this dense union is contained in (the preimage of) a Shimura subvariety of ${\mathcal{A}}_{3}$ illustrates the main theorem.

### v2

@@ 1. Introduction
Let $g\geq 1$ and let $C$ be a smooth connected complex projective curve of genus $g$. A point $x\in C(\mathbb{C})$ defines the Abel-Jacobi embedding $C\hookrightarrow J(C)$, where $J(C)$ is the Jacobian of $C$. Let $\iota\colon J(C)\rightarrow J(C)$ be the involution $y\mapsto-y$. Then the cycle $Z_{C}:=[C]-[\iota(C)]\in\mathrm{CH}^{g-1}(J(C))$ is a homologically trivial cycle on $J(C)$, and has a well defined image in $\mathrm{Griff}_{1}(J(C))$, the group of homologically trivial cycles modulo the group of algebraically trivial cycles (of dimension one). A famous theorem of Ceresa [Cer83] states that for $g\geq 3$, a very general smooth projective curve $C$ of genus $g$ has non-torsion Ceresa cycle in $\mathrm{Griff}_{1}(J(C))$.
@@
Let $g\geq 3$ and let $H_{g}$ denote the hyperelliptic locus in $\mathcal{M}_{g}$. There exists a finite set of subvarieties $T_{i}$ of dimension $\leq 2g-2$, and a finite set of preimages of strict Shimura subvarieties $X_{j}$ of $\mathcal{A}_{g}$, such that if $U$ is the complement of $(\bigcup_{i}T_{i})\cup(\bigcup_{j}X_{j})\cup H_{g}$ in $\mathcal{M}_{g}$, then the locus in $U(\mathbb{C})$ where $\nu$ is torsion is a countable set of points. Moreover, any positive dimensional component of the torsion locus of $\nu$ with Mumford-Tate group $\mathrm{GSp}_{2g}$ is defined over $\overline{\mathbb{Q}}$.
Hain [Hai24] and Gao–Zhang [GZ24, Theorem 1.3 (ii)] have, independently and with different methods, also constructed a Zariski dense open subset $U$ of $\mathcal{M}_{g}$ such that there exists at most countably many points in $U(\mathbb{C})$ where the Ceresa normal function is torsion.
@@
The locus of curves $C$ where $Z_{C}$ itself is torsion is contained in the torsion locus of $\nu$. In particular, by the standard spreading-out argument, the locus in $U$ of curves $C$ where $Z_{C}$ is torsion is a countable subset of $U(\overline{\mathbb{Q}})$. We note that a similar result have also been independently obtained by S-W. Zhang and Z. Gao [GZ24, Theorem 1.3-(i)].
@@
Theorem 1.3 has also been independently obtained by Gao–Zhang, see [GZ24, Corollary 1.9].
@@ 1.1. Ideas of the proofs
The proofs of the previous theorems rely on arithmetic and functional transcendence results in Hodge theory, namely the Manin–Mumford conjecture and recent results on the study of the Hodge locus for variations of mixed Hodge structures proved by Baldi–Urbanik [BU24]. The equidistribution part in Theorem 1.5 is inspired from the general equidistribution results established in [TT23] and relies on the computation of the generic rank of normal functions with big monodromy established in [Gao20a, Gao20b] and which relies on the Ax–Schanuel Theorem for variations of mixed Hodge structures.
@@ 1.2. Acknowledgement
The present work started during the session on open problems in the conference “The Ceresa Cycle in Arithmetic and Geometry” at ICERM. We thank the organizers for their efforts and ICERM for its hospitality, Hain for an inspiring talk, and Ziyang Gao, Philip Griffiths, Jef Laga, Aaron Landesman, Ari Shnidman and Shou-Wu Zhang for valuable discussions and comments. M.K. was supported by NSF grant DMS-2101482, and S.T. was supported by NSF grant DMS-2302388.
@@ 3.2. Analytic density and equidistribution of the torsion locus
In our setting, the rank of the Betti map was fully studied in [Gao20a], and a key ingredient was to use the mixed Ax-Schanuel Theorem. The following result is a particular case of [Gao20a, Theorem 1.4. (ii)]. Using this idea, we reproduce this result in a Hodge theoretic setting to make our paper more self-contained. The proof aligns more closely with [Gao20b, Theorem 9.1]. We refer also to [GZ24, Theorem 3.1] for a more general result in weight $-1$ and to [KU23] for a purely Hodge theoretic approach.
@@ 4.1. The torsion locus for the Ceresa cycle
We know from [LS24] that in genus 3, there is a ball quotient $Y\subset\mathcal{A}_{3}$ ruled by rational curves, countably many of which are contained in $\Phi(\mathrm{TL}_{\text{pos}}(\nu))$ for the Ceresa normal function. The torsion orders here are unbounded and the torsion locus is Zariski dense, as predicted by Theorem 3.9 and Remark 3.10 (as $\nu|_{Y}$ takes values in the Jacobian of an isotrivial VHS of rank 2); while the fact that this dense union is contained in (the preimage of) a Shimura subvariety of ${\mathcal{A}}_{3}$ illustrates the main theorem.